OpenStudy (anonymous):

A circle has its center at the origin, and (5, -12) is a point on the circle. How long is the radius of the circle?

OpenStudy (raden):

use the distance formula of two points (x1,y1) and (x2,y2) : d =sqrt ((x1-x2)^2 + (y1-y2)^2))

OpenStudy (anonymous):

Now let's see. The equation of a circle is given by:\[\bf (x-a)^2+(x-b)^2=r^2\]Where \(\bf (a,b)\) is the centre and \(\bf r\) is the radius. We know two things about this circle; firstly its centre is at the origin, i.e. \(\bf (a,b)=(0,0)\) and we have the already given point \(\bf (5,-12)\). So let's plug in this given information:\[\bf (x-0)^2+(x-0)^2=r^2 \implies x^2+y^2=r^2\]Now plug in the given point for x and y and solve for r:\[\bf (5)^2+(-12)^2=r^2 \implies r=?...\]Can you solve for 'r'?

OpenStudy (anonymous):

@marsa

OpenStudy (anonymous):

13?

OpenStudy (anonymous):

12? or 5?

OpenStudy (anonymous):

@marsa Your first answer was correct.

OpenStudy (jhannybean):

dont just guess....solve for it. :)

OpenStudy (anonymous):



OpenStudy (jhannybean):

lol..